# Lieb-Schultz-Mattis Anomalies and Anomaly Matching

### 01 Apr 2026 v1

@@ 1 INTRODUCTION
Understanding complex quantum many-body systems is an important but notoriously challenging task, as most models of such systems cannot be analytically solved and numerical approaches are limitations in various ways. Therefore, Lieb-Schultz-Mattis-type (LSM) constraints become highly valuable [1], which allow one to get important information about a system without solving its Hamiltonian. The typical statement of an LSM constraint starts with certain symmetry conditions where the symmetries include some lattice symmetries, and concludes that all states satisfying these conditions must have some nontrivial patterns of correlation and entanglement, and all local and even some non-local Hamiltonians obeying these conditions cannot have a unique gapped ground state (under periodic boundary conditions). Such LSM constraints have wide applicability, and the difference between different versions of them lies in the dimensionality of the system, the type of the symmetries, whether the system is a spin system or fermionic system, etc. An incomplete list of LSM constraints, some of which will be reviewed below, can be found in Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
Recently, LSM constraints have been interpreted as ’t Hooft anomalies [27, 28, 29, 30, 31], so we will use “LSM constraints” and “LSM anomalies” interchangeably. The notion of ’t Hooft anomalies originates from quantum field theory (QFT), where one considers a relativistic QFT with a certain internal symmetry $G$ [32]. Having the symmetry means that the theory can be coupled to a classical, background $G$ gauge field. The anomaly arises when the theory with the background gauge field cannot be made gauge-invariant [^1] More precisely, the gauge non-invariance cannot be remedied by adding local counterterms.. Consequently, the gauge field cannot be made dynamical, although the original theory is a consistent theory itself. Thus ’t Hooft anomalies in QFTs can be defined as an obstruction to gauging a symmetry. A well-known example of ’t Hooft anomalies is arguably the chiral anomaly of the axial/chiral U(1) symmetry of a Dirac fermion in odd spatial dimensions [33, 34].
@@
In the past decades, the study of symmetry-protected topological (SPT) phases in condensed matter physics [35] has led to significant advances in the understanding of ’t Hooft anomalies, beyond their original formulation in relativistic QFTs. In particular, it has been understood that quantum anomalies can be directly defined in lattice systems using the symmetry actions, where the symmetries may be internal or involve spatial transformations [36, 37, 12, 38, 15]. Compared to the continuum version, the lattice setting poses additional challenges in understanding these anomalies. For example, it is unclear how to make sense of background gauge fields for spatial symmetries. A completely general definition, characterization and classification of “lattice anomalies” is still an active research area and we will not go into this direction (see Refs. [39, 40, 41, 42] for recent development).
@@
The goal of this review is to give an elementary introduction to LSM anomalies and the anomaly matching. In Sec. 2, we discuss in detail how the LSM constraints manifest themselves in quantum spin chains and demonstrate the anomaly matching. In Sec. 3, we discuss how the LSM constraints discussed in Sec. 2 extend to higher dimensions. In Sec. 4, we describe the systematic framework for anomaly matching, and sketch how it is applied to different example systems. In Sec. 5, we discuss various other generalizations of LSM constraints. In Sec. 6, we discuss crystalline SPTs that are useful for understanding the LSM anomalies.
@@ 2.1 The LSM Theorem for Spin Chains
with periodic boundary condition $\bm{S}_{j+L}=\bm{S}_{j}$. The dots include other interactions invariant under SO(3) spin rotations and lattice translations. A version of the LSM theorems is the following statement [1], which applies to any local Hamiltonian of the form in Eq. (1).
@@
We can now state a general version of the LSM theorem in 1D with an internal symmetry and the lattice translation symmetry. Suppose the internal symmetry group of the system is $G$. $G$ is “on-site”, which means that the unitary transformation $U(g)$ for $g\in G$ takes the following form: $U(g)=\prod_{j}U_{j}(g)$, where $U_{j}(g)$ acts on the $j$-th site and identity elsewhere. We assume that all the symmetry action is translation invariant, that is, all sites transform in the same way.
@@
A projective representation is when the phase factors $\omega(g,h)$ cannot be “gauged away”, meaning that it cannot be written as $u(g)u(h)/u(gh)$ for some phase factors $u$ (otherwise one can redefine $U_{j}(g)\rightarrow u(g)U_{j}(g)$ to eliminate $\omega$). A familiar example is the half-integer-spin representations of SO(3).
@@
Theorem 2.2.
A translation-invariant, $G$-symmetric local Hamiltonian with a nontrivial $\omega]$ per unit cell cannot have a non-degenerate, gapped ground state.
@@
We remark that translation is not the only lattice symmetry that can lead to a mixed anomaly with internal symmetries. The space group of a 1D chain is quite simple, which is generated by translation and reflection. Reflection can also lead to LSM anomaly. To see it, we first consider a site-centered reflection ${\cal R}_{s}$ and discuss an example of LSM theorem that only requires SO(3) and ${\cal R}_{s}$. For simplicity, still consider the spin chain Hamiltonian on a periodic chain of $L$ sites. The reflection transformation ${\cal R}_{s}$ centered at site $1$ maps $j$ to 2j2-j mod $L$. We now argue that as long as the Hamiltonian preserves both SO(3) and ${\cal R}_{s}$, the ground state cannot be non-degenerate.
@@
On the other hand, in a spin-1/2 lattice with a bond-centered reflection symmetry ${\cal R}_{b}:j\rightarrow 1-j$ mod $L$, one can easily construct a spin-singlet, $\mathcal{R}_{b}$-invariant SRE state by pairing the spins on site $2j$ and $2j+1$ into a singlet. So there is no LSM anomaly associated with bond-centered reflection and spin rotation. As a lesson, we see that the position of the degrees of freedom (DOF) can determine whether there is LSM constraint. How should one determine and characterize the LSM constraints in different lattice systems? An answer to this question is provided by lattice homotopy [4, 8], which will be reviewed in Sec. 3.
@@ 2.2 Rigorous Approaches
The original proof for the anti-ferromagnetic Heisenberg chain in Ref. [1] is fully rigorous. However, the physical argument presented in Sec. 2.1 for discrete symmetry groups relies on the unproven assumption that the bulk gap remains robust upon introducing a symmetry defect (this has been rigorously established for U(1) symmetry defect [44]).
@@
Here having long-range correlation means that the connected correlator does not decay to zero for two local operators that are far away, and entanglement area law means that the von Neumann entropy for a segment of the spin chain is upper bounded when the length of the segment increases. This theorem means that if a system has an LSM anomaly, then any symmetric state must have some nontrivial correlation and entanglement. In particular, it implies that any symmetric state must be long-range entangled (LRE). We remark that the states discussed here do not have to be the ground state of any natural Hamiltonian. Similar theorems were also proved in Refs. [7, 10].
What is the relation between Theorems 2.1, 2.2 and Theorem 2.4? In fact, Theorems 2.1 and 2.2 can be derived and generalized using Theorem 2.4 and various other theorems. It is known that if a gapped local Hamiltonian has a unique ground state, then this ground state does not have long-range correlation [47]. Moreover, this ground state satisfies the entanglement area law [48]. So under the symmetry condition in Theorem 2.4, a symmetric gapped local Hamiltonian cannot have a unique ground state, which proves Theorems 2.1 and 2.2.
Interestingly, the locality condition in Theorems 2.1 and 2.2 can be relaxed. As proved in Refs. [13, 49], suppose a Hamiltonian is non-local but has interactions decaying faster than $1/r^{2}$, with $r$ the distance between the interacting DOF, if it is gapped and has a unique ground state, then its ground state satisfies the entanglement area law. Furthermore, Ref. [47] shows that such a ground state does not have long-range correlation. Therefore, we reach the following generalization of Theorems 2.1 and 2.2.
@@
Under the same symmetry condition as in Theorem 2.4, if a Hamiltonian is symmetric and its interaction decays faster than $1/r^{2}$, with $r$ the distance between the interacting DOF, then this Hamiltonian cannot have a unique gapped ground state.
We remark that the above way of thinking about the LSM theorem highlights the LSM anomaly as a constraint on the patterns of correlation and entanglement of the symmetric states, and the constraint on the spectra of symmetric Hamiltonians is its derived consequence. This perspective is powerful, because Theorem 2.4 only takes the symmetry condition as the input and makes no reference to any Hamiltonian, which underscores the kinematic nature of anomalies. Also, this approach leads to constraints on the spectra of non-local Hamiltonians under a wide class of symmetries, as stated in Theorem 2.5, which has not been obtained using other approaches (for the special case with $G_{\rm int}={\rm SO(3)}$, Theorem 2.5 was also proved in Ref. [14] by explicitly constructing low-lying states).
@@ 2.3 LSM Anomaly Matching in Low-Energy Theory
LSM anomaly by definition rules out completely trivial ground state. Its power however goes further than just the no-go theorem: The anomaly in fact puts strong constraints on the low-energy theory, known as “anomaly matching”. In this section, we demonstrate anomaly matching using the IR field theory of the spin-1/2 chain as an example.
@@
The two U(1) factors are generated by charges $Q_{m}$ and $Q_{w}$, whose eigenvalues are quantized to integers. Following the CFT terminology, the labels stand for momentum and winding, respectively. Suppose that the system has a length $L$ and periodic boundary condition, then
@@
More interestingly, one finds that the lattice translation becomes $e^{i\pi Q_{m}}e^{i\pi Q_{w}}e^{iPa}$. Here $P$ is the momentum operator of the field theory and $a$ the lattice spacing, so $e^{iPa}$ implements a continuous translation by $a$. Consequently, $e^{i\pi Q_{m}}e^{i\pi Q_{w}}$ is an exact symmetry in the low-energy theory, unlike the other elements of $\mathrm{U}(1)_{w}$. More precisely, $\mathrm{U}(1)_{w}$ is only emergent in the low-energy theory and is violated by irrelevant perturbations, except for its $\mathbb{Z}_{2}$ subgroup. Therefore, symmetries like $e^{i\pi Q_{m}}e^{i\pi Q_{w}}$ are termed “emanant symmetries” [29].
@@
Let us study the defect Hamiltonian $H[\alpha]$ using the IR theory. The continuum limit of Eq. (3) is  [^2] The same result can be reached by twisting the boundary condition of the fields by the symmetry $e^{i\alpha Q_{m}}$. From $e^{i\alpha Q_{m}}\theta e^{-i\alpha Q_{m}}=\theta+\alpha$, we impose the boundary condition $\theta(x+L)=\theta(x)+\alpha$ mod $2\pi$. As a result, the eigenvalues of $Q_{w}$ is modified to $N_{w}+\frac{\alpha}{2\pi}$, where $N_{w}\in\mathbb{Z}$ (while $Q_{m}$ still quantized to integers).
@@
What other field theories can “saturate” the anomaly? Let us limit ourselves to conformal field theories. Given the SO(3) symmetry, a natural class of CFTs to consider is the ${\rm SU}(2)_{k}$ WZW models. Using similar method, it is shown that $k$ must be odd (even) if ${\rm SU}(2)_{k}$ is realized as the IR theory of a half-integer (integer) spin chain [51, 52].
@@ 2.4 Filling Anomaly
The general LSM theorems for spin chains requires a nontrivial projective representation per unit cell. Another way to force a LSM-like constraint is to have “fractional charge” per unit cell. Again, for simplicity, consider a translation symmetric 1D system with U(1) charge conservation  [^3] We will assume that the total charge $Q$ is quantized to integers.. For example, this may be a spin-1/2 chain with $\mathrm{U}(1)$ spin rotation symmetry around a fixed axis, where the charge is the total magnetization. The filling fraction $\nu$ is defined as the average amount of charge per unit cell. We are interested in lattice systems with a fixed $\nu$. It is not difficult to generalize the argument in Sec. 2.1 to the present case. If we similarly define the defect Hamiltonian $H[\alpha]$, the modified translation operator $T(\alpha)$ satisfies
@@
Here $Q$ is the total charge. We can attempt to get rid of the phase factor in the first relation by redefining $T(\alpha)\rightarrow e^{i\alpha\frac{Q}{L}}T(\alpha)$, which changes it to $T(\alpha)^{L}=1$, but then the second relation is modified to $T(\alpha+2\pi)=e^{-2\pi i\nu}T(\alpha)$. In particular, as long as $\nu$ is not an integer, one finds obstruction to a non-degenerate gapped ground state similar to the one discussed in Sec. 2.1.

### 29 Apr 2026 v2

@@ 1 INTRODUCTION
Understanding complex quantum many-body systems is an important but notoriously challenging task, as most models of such systems cannot be analytically solved and numerical approaches are limited in various ways. Therefore, Lieb-Schultz-Mattis-type (LSM) constraints become highly valuable [1], which allow one to get important information about a system without solving its Hamiltonian. The typical statement of an LSM constraint starts with certain symmetry conditions where the symmetries include some lattice symmetries, and concludes that all states satisfying these conditions must have some nontrivial patterns of correlation and entanglement, and all local and even some non-local Hamiltonians obeying these conditions cannot have a unique gapped ground state (under periodic boundary conditions). Such LSM constraints have wide applicability, and the difference between different versions of them lies in the dimensionality of the system, the type of the symmetries, whether the system is a spin system or fermionic system, etc. An incomplete list of LSM constraints, some of which will be reviewed below, can be found in Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
Recently, LSM constraints have been interpreted as ’t Hooft anomalies [27, 28, 29, 30, 31, 32, 33, 34], so we will use “LSM constraints” and “LSM anomalies” interchangeably. The notion of ’t Hooft anomalies originates from quantum field theory (QFT), where one considers a relativistic QFT with a certain internal symmetry $G$ [35]. Having the symmetry means that the theory can be coupled to a classical, background $G$ gauge field. The anomaly arises when the theory with the background gauge field cannot be made gauge-invariant [^1] More precisely, the gauge non-invariance cannot be remedied by adding local counterterms.. Consequently, the gauge field cannot be made dynamical, although the original theory is a consistent theory itself. Thus ’t Hooft anomalies in QFTs can be defined as an obstruction to gauging a symmetry. A well-known example of ’t Hooft anomalies is the chiral anomaly of the axial/chiral U(1) symmetry of a Dirac fermion in odd spatial dimensions [36, 37].
@@
In the past decades, the study of symmetry-protected topological (SPT) phases in condensed matter physics [38] has led to significant advances in the understanding of ’t Hooft anomalies, beyond their original formulation in relativistic QFTs. In particular, it has been understood that quantum anomalies can be directly defined in lattice systems using the symmetry actions, where the symmetries may be internal or involve spatial transformations [39, 40, 12, 41, 15]. Compared to the continuum version, the lattice setting poses additional challenges in understanding these anomalies. For example, it is unclear how to make sense of background gauge fields for spatial symmetries. A completely general definition, characterization, and classification of “lattice anomalies” is still an active research area, and we will not pursue this direction here (see Refs. [42, 43, 44, 45] for recent developments).
@@
The goal of this review is to give an elementary introduction to LSM anomalies and anomaly matching. In Sec. 2, we discuss in detail how the LSM constraints manifest themselves in quantum spin chains and demonstrate the anomaly matching. In Sec. 3, we discuss how the LSM constraints discussed in Sec. 2 extend to higher dimensions. In Sec. 4, we describe the systematic framework for anomaly matching, and sketch how it is applied to different example systems. In Sec. 5, we discuss various other generalizations of LSM constraints. In Sec. 6, we discuss crystalline SPTs that are useful for understanding the LSM anomalies.
@@ 2.1 The LSM Theorem for Spin Chains
with periodic boundary conditions $\bm{S}_{j+L}=\bm{S}_{j}$. The dots include other interactions invariant under SO(3) spin rotations and lattice translations. A version of the LSM theorem is the following statement [1], which applies to any local Hamiltonian of the form in Eq. (1).
@@
We can now state a general version of the LSM theorem in 1D with an internal symmetry and the lattice translation symmetry. Suppose the internal symmetry group of the system is $G$. $G$ is “on-site”, which means that the unitary transformation $U(g)$ for $g\in G$ takes the following form: $U(g)=\prod_{j}U_{j}(g)$, where $U_{j}(g)$ acts on the $j$-th site and identity elsewhere. We assume that all the symmetry actions are translation invariant, that is, all sites transform in the same way.
@@
A representation is projective when the phase factors $\omega(g,h)$ cannot be “gauged away”, meaning that it cannot be written as $u(g)u(h)/u(gh)$ for some phase factors $u$ (otherwise one can redefine $U_{j}(g)\rightarrow u(g)U_{j}(g)$ to eliminate $\omega$). A familiar example is the half-integer-spin representations of SO(3).
@@
Theorem $2.$
A translation-invariant, $G$-symmetric local Hamiltonian with a nontrivial projective representation of $G$ per unit cell cannot have a non-degenerate, gapped ground state.
@@
We remark that translation is not the only lattice symmetry that can lead to a mixed anomaly with internal symmetries. The space group of a 1D chain is quite simple. It is generated by translation and reflection. Reflection can also lead to an LSM anomaly. To see it, we first consider a site-centered reflection ${\cal R}_{s}$ and discuss an example of the LSM theorem that only requires SO(3) and ${\cal R}_{s}$. For simplicity, still consider the spin chain Hamiltonian on a periodic chain of $L$ sites. The reflection transformation ${\cal R}_{s}$ centered at site 11 maps $j$ to 2j2-j mod $L$. We now argue that as long as the Hamiltonian preserves both SO(3) and ${\cal R}_{s}$, the ground state cannot be non-degenerate.
@@
On the other hand, in a spin-1/2 lattice with a bond-centered reflection symmetry ${\cal R}_{b}:j\rightarrow 1-j$ mod $L$, one can easily construct a spin-singlet, $\mathcal{R}_{b}$-invariant SRE state by pairing the spins on site $2j$ and $2j+1$ into a singlet. So there is no LSM anomaly associated with bond-centered reflection and spin rotation. As a lesson, we see that the position of the degrees of freedom (DOF) can determine whether there is an LSM constraint. How should one determine and characterize the LSM constraints in different lattice systems? An answer to this question is provided by lattice homotopy [4, 8], which will be reviewed in Sec. 3.
@@ 2.2 Rigorous Approaches
The original proof for the antiferromagnetic Heisenberg chain in Ref. [1] is fully rigorous. However, the physical argument presented in Sec. 2.1 for discrete symmetry groups relies on the unproven assumption that the bulk gap remains robust upon introducing a symmetry defect (this has been rigorously established for a U(1) symmetry defect [47]).
@@
Here having long-range correlation means that the connected correlator does not decay to zero for two local operators that are far away, and the entanglement area law means that the von Neumann entropy for a segment of the spin chain is upper bounded when the length of the segment increases. This theorem means that if a system has an LSM anomaly, then any symmetric state must have some nontrivial correlation and entanglement. In particular, it implies that any symmetric state must be long-range entangled (LRE). We remark that the states discussed here do not have to be the ground state of any natural Hamiltonian. Similar theorems were also proved in Refs. [7, 10].
What is the relation between Theorems 2.1, 2.2 and Theorem 2.4? In fact, Theorems 2.1 and 2.2 can be derived and generalized using Theorem 2.4 and various other theorems. It is known that if a gapped local Hamiltonian has a unique ground state, then this ground state does not have long-range correlation [50]. Moreover, this ground state satisfies the entanglement area law [51]. So under the symmetry condition in Theorem 2.4, a symmetric gapped local Hamiltonian cannot have a unique ground state, which proves Theorems 2.1 and $2.$
Interestingly, the locality condition in Theorems 2.1 and 2.2 can be relaxed. As proved in Refs. [13, 52], suppose a Hamiltonian is non-local but has interactions decaying faster than $1/r^{2}$, where $r$ is the distance between the interacting DOF. If it is gapped and has a unique ground state, then its ground state satisfies the entanglement area law. Furthermore, Ref. [50] shows that such a ground state does not have long-range correlation. Therefore, we reach the following generalization of Theorems 2.1 and $2.$
@@
Under the same symmetry condition as in Theorem 2.4, if a Hamiltonian is symmetric and its interactions decay faster than $1/r^{2}$, with $r$ the distance between the interacting DOF, then this Hamiltonian cannot have a unique gapped ground state.
We remark that the above way of thinking about the LSM theorem highlights the LSM anomaly as a constraint on the patterns of correlation and entanglement of the symmetric states, and the constraint on the spectra of symmetric Hamiltonians is its derived consequence. This is a powerful perspective, because Theorem 2.4 only takes the symmetry condition as the input and makes no reference to any Hamiltonian, which underscores the kinematic nature of anomalies. Also, this approach leads to constraints on the spectra of non-local Hamiltonians under a wide class of symmetries, as stated in Theorem 2.5, which has not been obtained using other approaches (for the special case with $G_{\rm int}={\rm SO(3)}$, Theorem 2.5 was also proved in Ref. [14] by explicitly constructing low-lying states).
@@ 2.3 LSM Anomaly Matching in Low-Energy Theory
An LSM anomaly, by definition, rules out a completely trivial ground state. Its power, however, goes further than just the no-go theorem: The anomaly in fact puts strong constraints on the low-energy theory, known as “anomaly matching”. In this section, we demonstrate anomaly matching using the IR field theory of the spin-1/2 chain as an example.
@@
The two U(1) factors are generated by charges $Q_{m}$ and $Q_{w}$, whose eigenvalues are quantized to integers. Following the CFT terminology, the labels stand for momentum and winding, respectively. Suppose that the system has length $L$ and periodic boundary conditions. Then
@@
More interestingly, one finds that the lattice translation becomes $e^{i\pi Q_{m}}e^{i\pi Q_{w}}e^{iPa}$. Here $P$ is the momentum operator of the field theory and $a$ is the lattice spacing, so $e^{iPa}$ implements a continuous translation by $a$. Consequently, $e^{i\pi Q_{m}}e^{i\pi Q_{w}}$ is an exact symmetry in the low-energy theory, unlike the other elements of $\mathrm{U}(1)_{w}$. More precisely, $\mathrm{U}(1)_{w}$ is only emergent in the low-energy theory and is violated by irrelevant perturbations, except for its $\mathbb{Z}_{2}$ subgroup. Therefore, symmetries like $e^{i\pi Q_{m}}e^{i\pi Q_{w}}$ are termed “emanant symmetries” [32].
@@
Let us study the defect Hamiltonian $H[\alpha]$ using the IR theory. The continuum limit of Eq. (3) is  [^2] The same result can be reached by twisting the boundary condition of the fields by the symmetry $e^{i\alpha Q_{m}}$. From $e^{i\alpha Q_{m}}\theta e^{-i\alpha Q_{m}}=\theta+\alpha$, we impose the boundary condition $\theta(x+L)=\theta(x)+\alpha$ mod $2\pi$. As a result, the eigenvalues of $Q_{w}$ are modified to $N_{w}+\frac{\alpha}{2\pi}$, where $N_{w}\in\mathbb{Z}$, while the eigenvalues of $Q_{m}$ are still quantized to integers.
@@
What other field theories can “saturate” the anomaly? Let us limit ourselves to conformal field theories. Given the SO(3) symmetry, a natural class of CFTs to consider is the ${\rm SU}(2)_{k}$ WZW models. Using a similar method, it is shown that $k$ must be odd (even) if ${\rm SU}(2)_{k}$ is realized as the IR theory of a half-integer (integer) spin chain [54, 55].
@@ 2.4 Filling Anomaly
The general LSM theorem for spin chains requires a nontrivial projective representation per unit cell. Another way to force an LSM-like constraint is to have “fractional charge” per unit cell. Again, for simplicity, consider a translation symmetric 1D system with U(1) charge conservation, and we assume that the total charge $Q$ is quantized to integers. For example, this may be a spin-1/2 chain with $\mathrm{U}(1)$ spin rotation symmetry around a fixed axis, where the charge $Q$ is defined as $Q=\sum_{i}(S_{i}^{z}-\frac{1}{2})$. The filling fraction $\nu$ is defined as the average amount of charge per unit cell. We are interested in lattice systems with a fixed $\nu$. It is not difficult to generalize the argument in Sec. 2.1 to the present case. If we similarly define the defect Hamiltonian $H[\alpha]$, the modified translation operator $T(\alpha)$ satisfies [32]
@@
Here $Q$ is the total charge. We can attempt to get rid of the phase factor in the first relation by redefining $T(\alpha)\rightarrow e^{i\alpha\frac{Q}{L}}T(\alpha)$, which changes it to $T(\alpha)^{L}=1$, but then the second relation is modified to $T(\alpha+2\pi)=e^{-2\pi i\nu}T(\alpha)$. In particular, as long as $\nu$ is not an integer, one finds an obstruction to a non-degenerate gapped ground state similar to the one discussed in Sec. 2.1.